### v6

# HaSPeR: An Image Repository for Hand Shadow Puppet Recognition

Abstract
Hand shadow puppetry, also known as shadowgraphy or ombromanie, is a form of theatrical art and storytelling where hand shadows are projected onto flat surfaces to create illusions of living creatures. The skilled performers create these silhouettes by hand positioning, finger movements, and dexterous gestures to resemble shadows of animals and objects. Due to the lack of practitioners and a seismic shift in people’s entertainment standards, this art form is on the verge of extinction. To facilitate its preservation and proliferate it to a wider audience, we introduce HaSPeR, a novel dataset consisting of 15,000 images of hand shadow puppets across 15 classes extracted from both professional and amateur hand shadow puppeteer clips. We provide a detailed statistical analysis of the dataset and employ a range of pretrained image classification models to establish baselines. Our findings show a substantial performance superiority of skip-connected convolutional models over attention-based transformer architectures. We also find that lightweight models, such as MobileNetV2, suited for mobile applications and embedded devices, perform comparatively well. We surmise that such low-latency architectures can be useful in developing ombromanie teaching tools, and we create a prototype application to explore this surmission. Keeping the best-performing model ResNet34 under the limelight, we conduct comprehensive feature-spatial, explainability, and error analyses to gain insights into its decision-making process. To the best of our knowledge, this is the first documented dataset and research endeavor to preserve this dying art for future generations, with computer vision approaches. Our code and data will be publicly available.
{IEEEImpStatement}
This research is an impetus towards utilizing AI tools to revitalize the hitherto underexplored cinematic art form of hand shadow puppetry. Such tools may help understand the creativity frontier in generative models, facilitate the development of applications to teach shadowgraphy, and unveil several prospects for entertainment. The existing works, though distally relevant to shadowgraphy, explore the digitization of such precursory art forms via approaches that have since been rendered primitive and obsolete. The novel dataset that we introduce in this paper, namely HaSPeR, consists of 15,000 diverse samples garnered from performance clips of variably skilled puppeteers. Our extensive benchmarking reveals that the task of classifying the puppet silhouettes is reasonably solvable using lightweight and convolutional feature extractor models, with accuracies of $94.97\%$ by ResNet34 and $92.38\%$ by MobileNetV2. HaSPeR, as a data resource for all intents and purposes, can be a potential stride towards systematically preserving this artistic practice.
{IEEEkeywords}
Hand shadow puppetry, art digitization, benchmark, computer vision, dataset curation, deep learning, image classification, mobile application, ombromanie, optical flow estimation, shadowgraphy, silhouette classification, transfer learning
1 Introduction
Ombromanie, the ancient art of hand shadow puppetry, is a form of art that involves the mesmerizing interplay of light and shadow through the construction and manipulation of shadow figures or silhouettes on a surface, typically a screen or a wall, using one’s hands, body, or props [1, 2]. The alias “cinema in silhouette”111 https://en.wikipedia.org/wiki/Shadowgraphy_(performing_art) is sometimes used to refer to this proto-cinematic medium of entertainment. Its working principle is very straightforward—the puppeteer adeptly positions their hands between a radiant light source and a translucent screen, consequently conjuring shadows and silhouettes that emulate different creatures, as shown in Figure 3.
Figure 1: Ombromanie in a nutshell. [^3] The shadowgraphy cliparts are adapted from ClipArt ETC, Florida Center for Instructional Technology, College of Education, University of South Florida. Link: https://etc.usf.edu/clipart/galleries/266-hand-shadow-puppetry
Despite its rich history and captivating allure across many cultures,444 https://www.geniimagazine.com/wiki/index.php/Shadowgraphy there exists a notable dearth of resources specifically tailored to this artistic domain. With properly annotated and sourced data, researchers could study the intricacies of hand silhouette movements, shapes, and storytelling techniques, thereby enabling the development of sophisticated Artificial Intelligence (AI) systems for automatic recognition, classification, or even generation of ombromanie performances [3]. The generation aspect is particularly relevant given the demonstrable impotency of AI image generator models in accurately creating hands and fingers [4]. Apart from that, the development of applications that can facilitate the learning of ombromanie has the potential to breathe new life into this waning art form [5]. In 2011, UNESCO recognized shadow puppetry as an endangered artistic tradition by adding it to the Intangible Cultural Heritage list [6], which is why it necessitates more preservatory apparatus and research efforts.
In tandem with this motivation, this work introduces a seminal addition to the realm of data resources, HaSPeR (Hand Shadow Puppet Image Repository), a methodically curated novel image dataset of hand shadow puppets. The dataset comprises an assemblage of 15,000 samples, that we painstakingly source and verify from 68 professional shadowgraphist clips and 90 amateur shadowgraphist clips. We label and categorize the images with utmost precision to elicit robustness in the image classification models that will undergo training with these images. The samples in HaSPeR are diverse in nature since the source clips are recorded in a plethora of different poses, orientations, and background lighting conditions of the translucent screen. We also inculcate silhouette motion diversity via optical flow estimation [7] in the frame extraction process. We conduct a detailed analysis of HaSPeR’s statistical characteristics. We also employ a variety of state-of-the-art (SOTA) pretrained image classification models to establish a performance benchmark for validating the integrity of the dataset. Additionally, we conduct a thorough evaluation of several facets of the ace ResNet34 model, including its feature representations, interpretability, explainability, and classification errors that it encounters. In an effort to assess the potential of digitized ombromanie teaching tools, we create a simple and lightweight prototype Android application using Flutter for classifying hand shadow puppet images from the phone’s camera feed. We posit that our dataset possesses the potential to offer a wealth of opportunities for exploration and analysis into the artistic domain of hand shadow puppetry.
Figure 2: A flowchart depicting the dataset construction process.
2 Literature Review
The recognition and classification of hand shadow puppet images are intriguing problemspaces in the context of deep learning, albeit relatively underexplored. After rigorously analyzing the existing pool of research on the topic, we could identify several quasi-related works.
2.1 Image Classification and Recognition
Among the pioneering endeavors in hand shadow image classification was that of Huang et al. [8], who created Shadow Vision—a system to emulate an immersive virtual shadow puppet theater experience, employing a user’s hand gestures over an overhead projector to control the creation and manipulation of objects within a 3D Open Inventor [^5] Open Inventor™ toolkit — https://www.openinventor.com/ environment. The chain of stages underlying the implementation of Shadow Vision were acquisition, segmentation, feature extraction, and recognition of the infrared shadow puppet images. They also adopted a 3-layer neural network and the centralized contour moments modeling technique, using 13 features (7 moments of the object, length, angle, and the 4 endpoints of the axis of inertia). The data used for this study isn’t publicly available, and the methodology can be deemed somewhat obsolete in the modern purview, due to being supplanted by the emergence of deep learning models. Some recent works explore different convolutional models to assess their efficacy in Indonesian shadow puppet recognition. Sudiatmika et al. [9], Sudiatmika and Dewi [10] used the deep CNN models, AlexNet [11] and VGG-16 [12], and constructed a dataset of 2,530 images spanning 6 classes of puppets from museums in Bali. They also experimented with other convolutional models, such as Mask R-CNN [13] and MobileNet [14], in two separate studies [15, 16].
In a similar spirit, our work is an endeavor towards establishing a performance benchmark of the recent SOTA feature extractor models for hand shadow puppet contour images, in a more large-scale and comprehensive manner.
2.2 3D Modeling and Human Motion Capture
One of the earliest works involving silhouettes is a study by Brand [17] that explored the mapping of monocular monochromatic 2D shadow image sequences of humans to animated 3D body poses, using a configural and dynamical manifold created from data with a topologically special hidden Markov model (HMM), acquired via the process of entropy minimization without resorting to any articulatory body model. Several advances in vision-based human motion capture and analysis since then have leveraged human silhouette templates [18, 19], more specifically, hand and finger silhouettes [20, 21, 22].
2.3 Robotics
Huang et al. [23] introduced computer vision-aided shadow puppetry with robotics by matching shape correspondences of input images. They claimed that due to the physical limitations of human arms, it is often not feasible to construct complex shadow forms. Instead, they developed a framework that enabled them to produce shadow images with the mechanical arms of a robot. The authors built a library of shadow images and used them to orient the robotic arms into a formation resembling the intended shadow puppet. The data used for this study isn’t publicly available.
2.4 Human-Computer Interaction
The authors of [24] proposed a framework for controlling two Chinese shadow puppets—a human model and an animal model, with the use of body gestures via a Microsoft Kinect sensor. Carr and Brown [25] conducted a similar work by building a real-time Indonesian shadow puppet storytelling application that is capable of mimicking the full-body actions of the user, using the Microsoft Kinect sensor. In order to leverage contactless gesture recognition (CGR) to teach traditional Chinese shadow puppetry to beginners, Tsai and Lee [26] developed a system using Leap Motion sensors. These studies on digitizing the art of shadow puppetry, or puppetry in general, were influenced to some extent by other similar works in the gesture recognition domain [27, 28, 6, 29, 30, 31]. Tang et al. [32] developed an intelligent shadow play system, called ShadowTouch, which includes a multidimensional somatosensory interaction module coupled with an automatic choreography module, to facilitate natural interaction between the shadow play figures and the human users.
The motif of our work tessellates well with the core objectives of the aforementioned research works. The utilization of digitized traditional arts serves as a means to preserve their inherent legacies, and HaSPeR can be a potent contribution to the contemporary pool of resources to facilitate such innovative digitization for ombromanie.
3 Dataset Construction
The series of steps involved in our data acquisition process is broadly divided into three tasks—(a) procuring the performance clips, (b) extraction of the frames, and (c) categorization of each sample frame with a proper label. Figure 2 portrays this workflow behind our dataset preparation. We incorporate manual oversight at each step of the dataset creation in order to reconcile any exigencies pertaining to the quality of HaSPeR.
Table 1: Statistical summary of HaSPeR.
  Silhouette Class Clips Sample Distribution
Pro. Nov.
Training
Validation
Total
  Bird 6 6 600 400 1000
Chicken 2 6 600 400 1000
Cow 2 6 600 400 1000
Crab 4 6 600 400 1000
Deer 6 6 600 400 1000
Dog 7 6 600 400 1000
Elephant 5 6 600 400 1000
Horse 8 6 600 400 1000
Llama 2 6 600 400 1000
Moose 3 6 600 400 1000
Panther 2 6 600 400 1000
Rabbit 4 6 600 400 1000
Snail 4 6 600 400 1000
Snake 3 6 600 400 1000
Swan 10 6 600 400 1000
  Total 68 90 9000 6000 15000
158
3.1 Collating Shadowgraphy Clips
At the outset of the process, we procure 68 different clips of 14 different professional shadowgraphists from YouTube. [^6] https://www.youtube.com The video sources are licensed under fair use and a list consisting of the links to all of them is available in our GitHub [^7] GitHub repository — https://github.com/Starscream-11813/HaSPeR repository. We record the relevant portions of the performance videos using the open-source recording software OBS Studio. [^8] Open Broadcaster Software®https://obsproject.com/ Six novice volunteer shadowgraphists collectively produce 90 additional clips, with each contributing one clip for every class. As a consequence, the total number of source clips aggregates to $68+(15\times 6)=158$.
Figure 3: Samples from each class of the dataset.
3.2 Extracting Samples
To mitigate the presence of excessively similar and redundant image samples, we extract frames from these clips at reasonable intervals of $k$ after downsampling the clips to a resolution of $1438\times 1080$. The values of $k$ are judiciously chosen for the clips of each class, and every $k$th frame is selected as a candidate image sample (e.g., with $k\approx 180$, $200$, $220$ for a 60 FPS clip). From this sequence of extracted candidate frames, we prioritize those exhibiting significant motion. To this end, we estimate the motion vector field by calculating the optical flow between the consecutive $t$th frame and the $(t+k)$th frame. The magnitude of this motion is quantified by the mean L2 norm of the resulting flow field. We retain the frame pairs with an average flow magnitude surpassing a certain requisite threshold $\tau$, thereby ensuring the inclusion of dynamically distinct frames. We synergistically amalgamate two optical flow estimation methods: the Lucas–Kanade (LK) method [33] and the Total Variation L1 Regularization (TV-$L^{1}$) method [34]. The undergirding assumption beneath the LK method is brightness constancy and spatial coherence of the flow in a local neighbourhood of the pixel (say, the patch $W$) under consideration. It employs a multi-scale gradient descent optimization approach for the constraint equation shown in Equation 1.
$I_{x}\cdot u+I_{y}\cdot v+I_{t}=0$ (1)
where, $I_{x}=\frac{\partial I}{\partial x}$ and $I_{y}=\frac{\partial I}{\partial y}$ are the spatial gradients of the image intensity $I$, and $I_{t}=\frac{\partial I}{\partial t}$ is the temporal gradient. $u$ and $v$ are the horizontal and vertical components of the optical flow vector, respectively. The motion is then estimated by iteratively minimizing the cost function in Equation 2 at increasingly granular image resolutions, from coarse to fine.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}=\operatorname*{arg\,min}_{u,v}\sum_{\{i,j\}\in W}\left[I_{x}(i,j)\cdot u_{ij}+I_{y}(i,j)\cdot v_{ij}+I_{t}(i,j)\right]^{2}$ (2)
which can be then solved with the closed-form solution in Equation 3.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}=\begin{bmatrix}\sum_{\{i,j\}\in W}I_{x}^{2}&\sum_{\{i,j\}\in W}I_{x}I_{y}\\ \sum_{\{i,j\}\in W}I_{x}I_{y}&\sum_{\{i,j\}\in W}I_{y}^{2}\end{bmatrix}^{-1}\begin{bmatrix}-\sum_{\{i,j\}\in W}I_{x}I_{t}\\ -\sum_{\{i,j\}\in W}I_{y}I_{t}\end{bmatrix}$ (3)
As evident from Figure 4(c), the LK method can track the edge movement of the homogeneous silhouette patches, given the slight texture offered by the penumbral region of the shadow. However, as it is limited by local window constraints, it fails to capture the global motion of the shadow puppet. To ameliorate this issue, we resort to the TV-$L^{1}$ method, which is a variational method that minimizes the total variation of the flow field, subject to the L1 norm of the data fidelity term, which together form the energy function $E$ in Equation 4.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{TV-}L^{1}}$ $=\operatorname*{arg\,min}_{u,v}E(u,v)$ (4)
[…]
$\begin{bmatrix}u\\ v\end{bmatrix}^{*}$ $=\operatorname*{arg\,max}_{u,v}\left(\left\lVert\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}\right\rVert_{2},\left\lVert\begin{bmatrix}u\\ v\end{bmatrix}_{\text{TV-}L^{1}}\right\rVert_{2}\right)$ (5)
Figure 4: Optical flow estimation of contiguous candidate frames from the ‘Horse’ class.
Then from this maximum combination flow field of resolution $M\times N$, we compute the average L2 norm $\overline{V}$ using Equation 6.
[…]
If $\overline{V}>\tau$, we retain the corresponding frame pair as candidate samples, otherwise we continue the process with the $(t+k)$th and the $(t+2k)$th frames. Table 1 encapsulates some essential statistical information related to HaSPeR and provides a superficial overview of the dataset.
3.3 Labeling
After the extraction of the frames, the samples undergo manual scrutiny by 3 annotators who are pursuing undergraduate studies in Computer Science and Engineering (CSE). If a series of contiguous samples prima facie exhibit substantial similarity, we only keep a single image from that set of samples. The rest are discarded to avoid redundancy and to instill diversity. Another criterion that dictates the legitimacy of an image sample is its intelligibility. If the majority of the annotators agree on the unintelligibility of a sample, they discard it in unison. After performing this omission of unsuitable samples for each class, we end up with 15 different directories of images, each containing the curated samples of a particular class. The images in these folders are then further partitioned into training and validation sets, maintaining a 60:40 split. We also pragmatically incorporate a proper distribution of the samples sourced from amateur clips over both the training and validation sets, to avoid making the latter unfairly difficult for the classification models.
4 Dataset Description
To provide a tangible exposition of the diverse samples in the dataset, Figure 3 presents a collection of representative images across all 15 classes. With minimally astute perspicacity, we can observe that the samples vary in terms of the nature of the backgrounds, the anatomical structure of the puppeteers’ hands, the photometric opacity and sharpness of the projected silhouettes, and a panoply of other aspects.
4.1 Background Variance
The hand shadow puppetry setup that a puppeteer’s crew arranges before the performance greatly dictates the nature of the background on which the shadow puppets are displayed. If the location of the light source is very near to the wall or the translucent screen, then we can observe an elliptical shadow contour on the background as evident in Figures 3(a) and 3(m). The angular directionality of the light also manifests a gradient effect on the background as can be seen in Figures 3(d) and 3(h). The temperature and color of the light emanated by the light sources onto the screens also add to the diversity. To achieve this, we use six different light sources—candlelight, an incandescent bulb, sunlight, a CFL light, an Epson EB-972 XGA projector, and an LED bulb—each with different color temperatures. Historically, many other light sources were used by shadowgraphists such as marrow-fat lamps, flame torches, halogen lamps, lime lights, etc. Figure 5 depicts an overarching illustration of the monocular polychromatic background lighting diversity that we maintain in HaSPeR. We avail the overhead projector to emit light from across the visible spectral range (380380380380750750750750 nm). Additionally, we use a random-patterned combination of these colors (colloquially referred to as the psychedelic pattern) as the backdrop for a subset of the organically created samples.
Figure 5: Light sources for background diversity in HaSPeR.
4.2 Nature of the Silhouettes
The positioning of the light source with respect to the puppeteer’s hands plays a role in shaping the shadows’ quality. As per the natural laws of optics, proximity to the light source yields crisp, well-defined shadows (e.g., Figure 6(a)), while increasing the distance fosters softer, more diffuse shadows (e.g., Figure 6(b)) with a central umbra and peripheral penumbra. The higher the contrast between the silhouettes and their respective backdrops, the more visible and well-contoured the shadow puppets are. The direction of the light source influences the orientation and shape of the shadows. Shadows cast by overhead lighting sources may appear elongated, while shadows cast by low-angle lighting sources may exhibit softer edges and less pronounced contrast and sharpness. Similarly, due to varied values of the lateral incident angle at which the light sources are kept relative to the screen’s normal, we see horizontally elongated and compressed shadows. The shadows also differ in terms of the magnitude of their opacity, i.e., the degree to which the hands prevent the transmission of light being projected onto the screen.
Figure 6: Samples with different silhouette properties.
4.3 Hand Anatomy and Stylistic Flair of the Puppeteers
The physiological properties of the puppeteers’ hands can vary significantly due to a combination of genetic factors, environmental influences, and lifestyle choices. These nuanced anatomical variations of the wrists, palms, and digits of the puppeteers, along with the different stylistic choices they employ in their choreography, contribute as yet another avenue of diversity of the image samples in HaSPeR. Human beings, by nature, exhibit morphometric variations in finger length, palm width, and forearm thickness based on age and gender. As such, the cohort of novice shadowgraphists ($n=6$) that we employ for the creation of amateur samples comprises a balanced gender representation of 3 males and 3 females, spanning an age range from 9 to 25 years. Among the adults, the hand anthropometric measurements are of $18.75\pm 1.55$ cm in length and $8.66\pm 0.77$ cm in width. For the minors, who obviously have proportionally smaller hand dimensions, these measurements are $14.23\pm 1.16$ cm and $6.73\pm 0.82$ cm respectively. The gender representation among the 14 professional shadow puppeteers is however imbalanced, with 12 male and 2 female shadowgraphists. Figure 7 pristinely demonstrates the morphological variations of hand shadow puppets belonging to the ‘Deer’ class due to anatomical and stylistic diversity.
Figure 7: ‘Deer’ samples with different artistic representations.
4.4 Comparative Analysis
4.4.1 Inter-class Similarity
Due to the conspicuous resemblance in the anatomical structures of certain animal species, the samples belonging to the classes corresponding to those animals exhibit a notable degree of similarity as well. Figures 3(e) and 3(j) are prime examples of such structural similitude that can be observed between the ‘Deer’ and ‘Moose’ classes. These similarities make the image classification task on HaSPeR quite challenging and culminate to being the reason behind a lot of misclassifications, as discussed in Section 5.3.
Figure 8: Classifier block attached to the tail-end of the pretrained models. Here, $d$ is the feature dimension of the anterior model. The output features of layer $l$ is $z^{[l]}=W^{[l]}a^{[l-1]}$, where $a^{[l-1]}$ denotes the activation values of the preceding $(l-1)$th layer. The batch normalized value of the $i$th output feature $z^{[l]}_{i}$ is $\hat{z}^{[l]}_{i}=\gamma^{[l]}z_{norm}^{[l](i)}+\beta^{[l]}$, where $\gamma$ and $\beta$ are learnable parameters. The activation values of layer $l$ are denoted by $a^{[l]}=g(\hat{z}^{[l]})$ which is computed using the activation function $g=\operatorname{ReLU}$. The predicted probabilities are determined by using the $\operatorname{softmax}$ function on the logits, i.e., $P(\hat{y_{i}}=1|x)=\operatorname{softmax}(y)_{i}=\frac{e^{y_{i}}}{\sum_{j=1}^{15}e^{y_{j}}}$
4.4.2 Intra-class Dissimilarity
Some classes include samples of multiple species of the same animal, and these samples are starkly different in appearance from one another. Given the presence of such quasi-disparate samples, along with the individualistic flair that manifests through the puppeteers’ stylistic choices, a particular class may show a lot of intra-class dissimilarity. As aforementioned, Figure 7 portrays the heterogeneity of this nature among the samples from the ‘Deer’ class.
4.5 Statistical Analysis
Table 1 presents the statistical properties of the HaSPeR dataset. It tabulates the proportion of samples belonging to each of the 15 classes and their corresponding training-validation splits. While searching for hand shadow puppetry performances, we anecdotally observe that certain classes of puppets are more popular than others. Irrespective of this fact, we make sure all 15 classes of puppets are equitably represented in our dataset, with each class having 1,000 image samples ($\approx 6.\dot{6}\%$). As evident in Table 1, the image samples are evenly distributed across all 15 classes. The proportions of professionally sourced samples belonging to the ‘Llama’ and ‘Snake’ classes ($14\%$ and $27.3\%$ respectively) are slightly low due to a scarcity of performance clips starring hand shadow puppets of these classes. For cases such as these, we supplement the classes with samples organically created by our novice shadowgraphist cohort. Each class has $\approx 47.827\pm 1.414\%$ samples from clips of professional performers, and the rest $\approx 52.172\pm 1.414\%$ samples are sourced from amateur clips. Of the 15,000 samples, approximately $76.53\%$ feature male hands, while $23.64\%$ represent female hands. In tandem, considering the target demographic of ombromanie, around $28.33\%$ of the samples in HaSPeR consist of children’s hands. In totality, we end up with 9,000 samples in the training set and 6,000 samples in the validation set, thereby partitioning HaSPeR by maintaining a 60:40 ratio.
5 Developing Benchmark for HaSPeR
A series of pretrained models are used as feature extractors to develop a benchmark for the dataset. The models are pretrained on the ImageNet [35] dataset and fine-tuned on HaSPeR. We implement the training pipeline using the Pytorch [^9] https://pytorch.org/vision/stable/models.html framework. This section presents an overview of the models, evaluation metrics, and experimental results.
5.1 Experimental Setup
5.1.1 Baseline Models
For this classification task, we use 31 feature extractor models as baselines, which are listed in Table 2. Some of these models have a track record of good performance across various other image classification tasks [36]. We examine both conventional Convolutional Neural Networks (CNNs) and CNNs augmented with attention mechanisms. Some models have multiple variants in terms of size or number of parameters, and we compare the performance among those variants as well.
5.1.2 Performance Metrics
We use top-$k$ validation accuracy values (with $k=1,2,3$), Precision, Recall, and F1-score as evaluation metrics to perform comparative analyses among the aforementioned models. The latter three judgment criteria are used due to the slightly imbalanced nature of HaSPeR, as evident in Table 1.
5.1.3 Classifier Network
We adopt two approaches to arrive at the final 15-dimensional layer since there are a total of 15 classes to predict from. The first approach is to directly append a 15-dimensional fully connected layer at the tail-end of the vanilla models. The second approach incorporates the classifier block portrayed in Figure 8.
5.1.4 Hyperparameters and Optimizer
We use Stochastic Gradient Descent (SGD) [37], with a learning rate $\alpha=0.001$ and momentum $\gamma=0.9$, as the optimizing method, and Cross Entropy Loss as the loss metric for all the models. To decay the learning rate, we use Step Scheduler, which decays $\alpha$ by $0.1$ every 5 epochs. Each model undergoes training for 50 epochs to ensure equitable comparison, and we empirically ascertain that 50 epochs are sufficient for all of the models to achieve convergence.
Table 2: Performance comparison of the vanilla and modified versions of the image classification models.
  Models Params. Performance Metrics
Vanilla w/ Classifier Block
Top-$k$ Accuracy (%) Precision Recall F1-score Top-$k$ Accuracy (%) Precision Recall F1-score
Top-$1$ Top-$2$ Top-$3$ Top-$1$ Top-$2$ Top-$3$
                ShuffleNetV2X10 [38] 2.3M 61.73 78.41 86.1 0.6559 0.6173 0.597 88.73 93.98 96.1 0.8995 0.8873 0.8853
ViTB16 [39] 86.6M 69.71 77.6 83.28 0.7276 0.6972 0.6969 68.88 76.65 81.36 0.7192 0.6868 0.6851
ViTL32 [39] 306.5M 85.1 91.56 94.48 0.872 0.851 0.8509 84.71 91.8 94.08 0.8632 0.8472 0.8465
AlexNet [11] 61.1M 87.01 93.61 95.46 0.884 0.8702 0.8708 88.18 92.58 94.8 0.8887 0.8818 0.8809
SqueezeNet1_1 [40] 1.2M 87.56 92.45 94.15 0.888 0.8757 0.8744 86.21 92.48 94.65 0.8754 0.8622 0.8637
MobileNetV3Small [41] 2.5M 89.48 94.31 95.76 0.9038 0.8948 0.8942 89.85 94.35 96.48 0.9082 0.8985 0.8976
SwinB [42] 87.8M 90.5 95.38 97.4 0.9128 0.905 0.9042 90.2 95.4 97.08 0.9097 0.902 0.9006
GoogLeNet [43] 6.6M 90.73 94.65 95.7 0.9105 0.9073 0.9059 92.18 95.65 96.58 0.9283 0.9218 0.9206
ResNet18 [44] 11.7M 90.91 95.28 96.6 0.9176 0.9092 0.9069 91.25 95.43 97.05 0.9229 0.9125 0.9119
MobileNetV3Large [41] 5.5M 91.2 94.48 95.98 0.9185 0.912 0.911 90.4 94.53 95.26 0.9147 0.904 0.9024
ConvNeXt [45] 88.6M 91.46 96.33 98.05 0.922 0.9147 0.914 92.55 96.36 97.96 0.9306 0.9255 0.9246
SwinV2B [46] 87.9M 91.58 96.25 97.61 0.921 0.9158 0.9151 91.48 96 97.55 0.9209 0.9148 0.9144
VGG16 [12] 138.4M 91.61 95.08 96.65 0.9248 0.9162 0.9168 91 95.21 96.45 0.9235 0.91 0.9119
MNasNet13 [47] 6.3M 91.66 95.65 97.01 0.924 0.9167 0.9149 91.45 95.86 97.26 0.9231 0.9145 0.9133
ConvNextLarge [45] 197.8M 91.88 95.9 97.7 0.9254 0.9188 0.9181 88 94.7 96.56 0.8942 0.88 0.8782
EfficientNetB0 [48] 5.3M 91.93 95.26 96.71 0.9257 0.9193 0.9178 90.4 93.75 95.1 0.9131 0.904 0.9022
MaxViT [49] 30.9M 92.01 96.5 97.81 0.9268 0.9202 0.9214 92.08 95.98 97.36 0.932 0.9208 0.9237
EfficientNetV2S [50] 21.5M 92.31 95.75 96.76 0.9375 0.9232 0.9245 94.45 97.35 98.3 0.9498 0.9445 0.9438
VGG19 [12] 143.7M 92.36 95.13 96.1 0.9354 0.9237 0.9242 91.8 95.06 96.15 0.9296 0.918 0.9187
MobileNetV2 [51] 3.5M 92.38 94.98 96.05 0.9303 0.9238 0.9233 92.31 95.38 96.91 0.9311 0.9232 0.9225
WideResNet50_2 [52] 68.9M 92.46 96.28 97.28 0.9331 0.9247 0.9235 93.35 95.73 97.15 0.9421 0.9335 0.933
ResNet50 [44] 25.6M 92.58 95.56 96.75 0.9332 0.9258 0.9252 93.08 96.48 97.2 0.9363 0.9308 0.9299
RegNetX32GF [53] 107.8M 92.86 95.71 96.93 0.9348 0.9287 0.9269 92.91 95.71 96.95 0.9366 0.9292 0.9282
DenseNet121 [54] 8.0M 92.93 95.75 96.88 0.9367 0.9293 0.9282 92.95 95.51 96.56 0.936 0.9295 0.9285
ResNext101_32X8D [55] 88.8M 93 96.41 97.23 0.9364 0.931 0.9303 94.2 96.61 97.58 0.952 0.942 0.9423
WideResNet101_2 [52] 126.9M 93.36 95.81 96.9 0.9423 0.9337 0.9332 92.73 96.35 97.63 0.9337 0.9273 0.9267
InceptionV3 [56] 27.2M 93.5 96.48 97.35 0.9401 0.935 0.9338 93.71 96.36 97.06 0.9446 0.9372 0.9371
DenseNet201 [54] 20.0M 93.56 95.78 96.73 0.945 0.9357 0.9353 94.43 97 97.61 0.9492 0.9443 0.9442
ResNet101 [44] 44.5M 93.81 96.23 97.71 0.9432 0.9382 0.9406 93.23 96.93 98.13 0.9386 0.9323 0.9321
ResNet152 [44] 60.2M 94.06 97.06 98.05 0.9447 0.9407 0.9394 93.05 96.73 97.48 0.9374 0.9305 0.9297
ResNet34 [44] 21.8M 94.97 97.23 98.23 0.9516 0.9497 0.9491 91.98 95.95 97.2 0.9266 0.9198 0.9189
5.1.5 Data Augmentation and Preprocessing
In order to generate a more diverse pool of training samples, we also incorporate data transformation techniques [^10] https://pytorch.org/vision/stable/transforms.html—Random Resize, Random Perspective, Color Jitter, Random Invert, Random Horizontal Flip, Random Crop, Random Rotation, Gaussian Blur, and Random Affine with translation and shearing—while training the models. We choose these data augmentation techniques since the classes in HaSPeR are mostly rotationally asymmetric and incongruent. Consequently, the augmented samples aid in eliciting better generalization abilities and robustness for all the models. The input images that are fed to the models are appropriately resized a priori using Bicubic Interpolation.
5.2 Results and Findings
5.2.1 Performance Analysis
The ResNet34 model yields the best performance with a top-$1$ accuracy of $94.97\%$. The vanilla version of the model also yields the highest top-$2$ accuracy, top-$3$ accuracy, Precision, Recall, and F1-scores of $97.23\%$, $98.23\%$, $0.9516$, $0.9497$, and $0.9491$ respectively. Upon being equipped with the classifier block shown in Figure 8, the EfficientNetV2S model yields the highest top-$k$ accuracies and Recall. In contrast, the ResNext101_32X8D and DenseNet201 models demonstrate the best performance across Precision and F1-score metrics respectively. At this recess of the performance analysis, we consider the top-$1$ accuracy metric to be the most statistically significant metric. As evident in Table 2, the vanilla models listed in the penultimate row and above lag behind the ResNet34 model when it comes to the top-$1$ accuracy value (as well as the other metrics), which is why we adjudicate that ResNet34 is the best-performing model.
Figure 9: $t$-Distributed Stochastic Neighbor Embedding ($t$-SNE) feature representations of ResNet34.
5.2.2 Feature Space Visualization and Analysis
In order to visualize the learned feature space of ResNet34, we resort to the dimensionality reduction technique called $t$-Distributed Stochastic Neighbor Embedding [57] since it can preserve the proximity of high-dimensional data points. For high-dimensional data residing on or proximate to a low-dimensional, non-linear manifold, it becomes imperative to preserve this proximity of the collapsed low-dimensional representations for closely resembling data points. Achieving such proximity preservation is often unattainable through linear mappings such as Principal Component Analysis (PCA), which is why we opt for the $t$-SNE dimensionality reduction approach. We can pragmatically infer from the 2D-collapsed visualizations of the high-dimensional feature representations in Figure 9, that the classes are nicely clustered and congealed with minimal overlaps and outliers. This enables the model to easily determine the decision surface in the high-dimensional feature space and perform very well on the classification task.
5.2.3 Qualitative Analysis and Explainability
As depicted in Figure 11, we adopt a plethora of explainable AI (xAI) techniques for the best-performing ResNet34 model to understand its decision-making. While viewing the GradCAM (Gradient-weighted Class Activation Mapping) [58] attention heatmaps, it becomes apparent that the model puts more gravitas on the common-sense distinguishing traits. For example, in Figure 11(b), we observe the regions of the image samples predominantly influencing their respective classification scores—the wingspan and beak of a bird, the gallinaceous comb of a chicken, the horns and concave head of a cow, the appendages of a crab, the horns of a deer, the long-slanted head of a dog, the tusks of an elephant, the long maxilla-mandibular jaw of a horse, the long-eared and tapered head of a llama, the upright horns of a moose, the big eyes and small ears of a panther, the petite hands and head of a rabbit, the shell and antennae of a snail, the lateral hood expansion of a snake, as well as the slender neck and wing feathers of a swan. As human beings, we evoke these same distinguishing characteristics while classifying the images using our own visual reasoning faculties. As exemplified in Figure 11(c), for local interpretation, we use the model-agnostic technique called LIME (Local Interpretable Model-agnostic Explanations) [59]. The green highlights indicate regions of the image that contribute positively to the probability of the assigned label, while the red highlights signify areas that reduce this probability. We also demonstrate the spatial support of the top-$1$ predicted classes by generating the saliency maps [60] in Figure 11(d). These maps are rendered using a solitary back-propagation pass through the ResNet34 model, and they accentuate the salient areas of the given image, characterized by their discriminative attributes with respect to the given class.
Figure 10: Android application for shadow puppet recognition.
Figure 11: The juxtaposition of original image samples from the HaSPeR dataset with their corresponding GradCAM Heatmaps, LIME Visualizations, and Saliency Maps (for the best-performing vanilla ResNet34 model).
5.2.4 Practicality Analysis as a Teaching Tool
It is noteworthy to point out that MobileNetV2, with only 3.5 million parameters, managed to surpass many of the other models in terms of performance. This indicates the suitability of this image classification task for lighter, low-latency models that can be used in mobile applications and embedded devices. We create a simple prototype Android application using Flutter to test the efficacy of MobileNetV2 in classifying hand shadow puppet images from the phone’s camera feed. In order to make the prototype work as seamlessly as intended, we make sure the vicinity is well-lit, and the camera accurately captures a sharply focused silhouette. We find that the model has a memory footprint of $29$ MB and achieves an average inference time of $880$ $\mu$s on the Snapdragon 8 Gen 2 mobile chipset featured in the Samsung Galaxy S23 smartphone. Figure 10 portrays the snapshots of the prototype application. There are several other practical implementation challenges involved in this endeavor that we can identify for an educational mobile application to comport well with the target demographics and real-world settings. The system must operate with minimal computational overhead, ensure real-time responsiveness, and maintain low latency. This necessitates a pragmatic tradeoff between the FLOP count and classification accuracy, with the requisite model compression and optimization techniques. Given the potential variability in device camera capabilities, the application must have preprocessing steps including, but not limited to, denoising, adaptive contrast enhancement, and sharpening the input feed to mitigate artifacts. In recognition of the diverse motor capabilities of users, particularly younger learners and users with dexterity impairments, the application must have intelligent motion compensation to stabilize the shaky camera inputs. Adding to the desiderata is the consideration of a suitable UI/UX that is tailored for the pediatric user base, as is done in handwriting teaching apps for children [61], because an intuitive and enjoyable learning experience is of paramount importance for educational apps. This includes providing step-by-step tutorials, interactive guides with progressive difficulty, and illustrative diagrams to demonstrate the creation of hand shadow puppets. A holistically sound approach to this application development endeavor will surely potentiate the pedagogical feasibility of ombromanie.
Figure 12: Confusion Matrix of vanilla ResNet34.
Figure 13: Misclassified samples with visually similar samples of the predicted class.
5.3 Error Analysis
The confusion matrix for the ResNet34 model on our dataset, presented in Figure 12, reveals that the ‘Crab’ class exhibits the highest count of misclassifications. One obvious reason for this is the somewhat significant inter-class similarity among the ‘Bird’, ‘Moose’, ‘Rabbit’, and ‘Crab’ classes. Most of the misclassified samples are from visually similar classes. We can posit that navigating the intricacies of visually similar classes poses a significant challenge in this image classification task, as evident from the other pale-red entries of the confusion matrix in Figure 12. Even to the keen human eye, distinguishing between these classes may be perplexing, as they share common visual features, shapes, or color patterns that result in a high degree of resemblance. We examine various aspects, such as the distinctive features or characteristics that might have led to confusion and the degree of similarity between the misclassified classes. Figure 13(a) and 13(f) show the confusion between a ‘Crab’ sample and a ‘Rabbit’ sample which look visually quite similar. The same holds for one of the six ‘Moose’ samples that are misclassified as ‘Bird’ samples by the ResNet34 model, as depicted in Figure 13(b) and 13(g). We observe that misclassifications of this type occur when images belonging to different, but visually akin categories, are erroneously assigned to the wrong class. Another reason for misclassifications is the ambiguity of shape present in mid-action frames. For instance, the ‘Bird’ sample in Figure 13(h) is a transition frame between two successive wing flaps. However, due to the presence of this ambiguous sample in the training set of the ‘Bird’ class, the ResNet34 ends up misclassifying the ‘Panther’ sample in Figure 13(c) as a ‘Bird’ sample. The misclassification portrayed by the pair of Figure 13(d) and 13(i) is due to the combination of poor lighting and ineptitude of the amateur child puppeteer in creating ‘Llama’ shadows. The model confuses the ear protrusions of the ‘Llama’ sample to be the tentacular eyes of a ‘Snail’ sample. For Figure 13(e), we can pontificate the misclassification reasons to be the overlapping of the fingers and the distorted angle at which the sample was captured. The ‘Snail’ sample thereby gets wrongly classified as a ‘Bird’ sample due to the existence of the analogous sample portrayed in Figure 13(j). The green entries along the diagonal of the confusion matrix in Figure 12 indicate the reasonably good classwise prediction performance of the ResNet34 model, which is, to some extent, due to the perfectly balanced sample distribution in HaSPeR.
5.4 Avenues of Improvement
To reduce the number of misclassifications, models need to be imbued with the ability to learn certain nuanced features. In light of the contemporary image classification literature, we can opt to use Kolmogorov–Arnold Networks (KAN) [62] instead of a simple Multilayer Perceptron (MLP) as the classifier block. The use of Convolutional KANs [63] has yielded good results in many image classification benchmarks. The task of image classification on HaSPeR can be dubbed as a silhouette classification task, which is why we can leverage topological features of the shadow contours to achieve better results [64]. The silhouette polygonization algorithm (PoG), as proposed by Göçmen and Akata [65], may aid in achieving better classification accuracy. Other possible avenues may involve the use of ensemble methods coupled with voting schemes, or resorting to data augmentation with synthetically generated samples, but we defer the exploration of these hypotheses for future research endeavors.
6 Conclusion and Future Work
In this paper, we explore an intriguing subject matter, hand shadow puppetry, in the realm of computer vision. We introduce HaSPeR, a novel dataset with a sizable collection of 15,000 hand shadow puppet images distributed across 15 classes, taken from both expert and amateur puppetry performance clips, by dint of a sophisticated frame extraction process involving optical flow estimation. We fine-tune 31 pretrained image classification models on HaSPeR to establish a benchmark for the dataset. To explain and analyze the performance of the most erudite model, ResNet34, in comparison with other baseline models, we visually manifest its feature space using the $t$-SNE dimensionality reduction technique. We also perform thorough qualitative and error analyses for the ResNet34 model. We envisage the possibility of developing applications for imparting the art of shadowgraphy, via mobile and embedded devices. We claim that this work is novel and significant since it is the first publicly available dataset and study on image classification benchmarking that focuses only on ombromanie. There are many avenues in our work that warrant further investigation. We hope to reconcile those desiderata by enriching our dataset with numerous permutations of arm positions and finger movements, preferably by employing more skilled individuals with varying palm and wrist structures, thereby creating more diverse silhouettes. We also plan to experiment with a gesture detection technology such as MediaPipe [^11] MediaPipe — https://developers.google.com/mediapipe or Microsoft Kinect [^12] Kinect for Windows — https://learn.microsoft.com/en-us/windows/apps/design/devices/kinect-for-windows for leveraging depth coordinates of hand landmarks [66], and assess their efficacy in classifying hand shadow puppets.
7 Acknowledgments
We convey our heartfelt gratitude, in advance, to the anonymous reviewers for their constructive criticisms and insightful feedback which will surely be conducive to the improvement of the research work outlined in this paper. We also appreciate the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT) for the generous provision of computing resources during the course of this project. Additionally, we wish to acknowledge Shahriar Ivan, Department of Computer Science and Engineering, Islamic University of Technology, for his invaluable assistance in proofreading and offering a preliminary review of this manuscript. We further thank Mohammad Ishrak Abedin and Reaz Hassan Joader, from the same department, for lending their assistance, which was instrumental in refining this paper’s illustrations and diagrams. Syed Rifat Raiyan, in particular, wants to thank his parents, Syed Sirajul Islam and Kazi Shahana Begum, for everything.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The codes and datasets generated, written, and/or analyzed during the research project are available in the HaSPeR GitHub repository (https://github.com/Starscream-11813/HaSPeR) and in the HaSPeR Hugging Face repository (https://huggingface.co/datasets/Starscream-11813/HaSPeR).
Funding Sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. All associated costs were self-funded by the authors. Access to computing resources was provided by the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT).